-6(r+2)+4r-3=2
9(q+3)+2q-9=9
2(2w+3)-3w+7=10
-2(3h+2)-4h-6=5
-7(-2y+3)+4y-3=3
Need help with these 5 problems on transitional math.

Answer by macston(5194)

.
1).
-6(r+2)+4r-3=2
-6r-12+4r-3=2
-2r-15=2
-2r=17
r=-17/2
2).
9(q+3)+2q-9=9
9q+27+2q-9=9
11q+18=9
11q=-9
q=-9/11
3).
2(2w+3)-3w+7=10
4w+6-3w+7=10
w+13=10
w=-3
4).
-2(3h+2)-4h-6=5
-6h-4-4h-6=5
-10h-10=5
-10h=15
h=-15/10=-3/2
5)
-7(-2y+3)+4y-3=3
14y-21+4y-3=3
18y-24=3
18y=27
y=27/18=3/2